OpenStudy (anonymous):

Generate the first 5 terms of this sequence: f(1) = 2 and f(2) = 3, f(n) = f(n - 1) + f(n - 2), for n > 2. 2, 3, 5, 7, 9 2, 3, 4, 7, 11 2, 3, 5, 9, 11 2, 3, 5, 8, 13

OpenStudy (anonymous):

help please

OpenStudy (anonymous):

f(5)=f(5-1)+f(5-2) =f(4) + f(3) because 5-1 is 4 and 5-2 is 3 =2 + 1 because f(4) is 2 and f(3) is 1 =3

OpenStudy (adamaero):

try values for n that make is so it comes out to f(1) and f(2)

OpenStudy (anonymous):

The answer SHOULD be B....

OpenStudy (anonymous):

f(3) = 12 f(1) = 2 and f(2) = 5

OpenStudy (anonymous):

thanks cutie can we talk

OpenStudy (anonymous):

Surree :) and no problem

OpenStudy (anonymous):

im on my email